OpenStudy (anonymous):

Write and solve the proportion: 3 is to 12 as w is to 8.

OpenStudy (anonymous):

2

OpenStudy (anonymous):

3/12 = 2/8

OpenStudy (anonymous):

3/12 = w/ 8 cross multiply to get 24 = 12w solve for w to get 2

OpenStudy (anonymous):

3/12=w/12 w=3*8/12=2